Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

die

Expert replies
by ruplun » Fri Jul 29, 2011 5:18 am
A die with x sides has consecutive integers on its sides.If the probability of not getting a 4 on either of the 2 tosses is 36/49, how many sides does the die have?

a. 4
b. 5
c. 7
d. 8
e. 13
Join the discussion
Source: — Problem Solving |

by naveen451 » Fri Jul 29, 2011 9:02 am
ruplun wrote:A die with x sides has consecutive integers on its sides.If the probability of not getting a 4 on either of the 2 tosses is 36/49, how many sides does the die have?

a. 4
b. 5
c. 7
d. 8
e. 13
IMO C

I AM NOT SURE BUT
explanation

if suppose there are 7 sides
then probability of getting 4 =1/7

probability of not getting 4=1-1/7 =6/7

probability pf not getting 4 for two times =6/7*6/7=36/49

thus die has 7 sides.

if i am wrong plz correct me
Join the discussion

by tauseef10jan2004 » Wed Aug 03, 2011 2:24 am
let the die has n sides

then probability of getting 4 = 1/n
probability of not getting 4 = 1-(1/n) = (n-1)/n

probability of not getting 4 twice = ((n-1)/n)*((n-1)/n) = sqr((n-1)/n) = 36/49

i.e. (n-1)/n = 6/7
on solving this n = 7

answer is c. 7
Join the discussion

by edge » Wed Aug 03, 2011 3:58 am
I just saw 7^2 in the denominator and chose C, knowing that in most of the probability-of-die-rolls questions I have faced, the solution often had 6^n in the denominator (n = the number of rolls of a standard dice).
Join the discussion

by shashank.ism » Wed Aug 03, 2011 4:33 am
Its quite simple. Let there be n sides of die.
so probability of getting 4 in one throw = 1/n
and probability of not getting 4 in one throw = (n-1)/n
so probability of not getting a 4 i n both the 2 throw = {(n-1)/n}^2
--> {(n-1)/n}^2 = 36/49
--> {(n-1)/n} = 6/7
--> n = 7.
Hence Ans C
My Websites:
www.mba.webmaggu.com - India's social Network for MBA Aspirants

www.deal.webmaggu.com -India's online discount, coupon, free stuff informer.

www.dictionary.webmaggu.com - A compact free online dictionary with images.

Nothing is Impossible, even Impossible says I'm possible.
Join the discussion